# neutron star; dark matter; gravitational waves; tidal deformability; neutron decay

### v2

@@ 3.5 Speed of Sound in Dark Matter Admixed Neutron Stars
The stiffness of matter in compact objects is described by the sound speed $c_{s}$, which plays a critical role in preventing a static NS from collapsing into a BH. Unlike the mass or the radius, the sound speed cannot be directly measured by experiments, nor it is easily inferred. However, the tidal deformability, $c_{s}$, is still a key parameter for characterizing the NS properties and, in particular, for quantifying how resistant the matter is to compression in extreme conditions. In this regard, Ecker et al. Ecker and Rezzolla (2022) have demonstrated that in light stars (whose $c_{s}$ increases gradually towards the center), the outer layers are soft while the core is rather stiff; whereas heavy NSs have a stiffer outer shell and a softer core (where the sound speed reaches a local minimum).
@@ 5 Neutron Decay to Dark Matter
DM interactions can be constrained by investigating neutron decay within NSs. The longstanding discrepancy in neutron decay lifetime, observed using the beam and bottle methods Berezhiani and Bento (2006); Serebrov et al. (2008); Fornal and Grinstein (2018), provides an avenue to explore these interactions. In the bottle method, neutrons are trapped, making it sensitive to the total neutron width, yielding a slightly shorter lifetime of $\tau_{n}^{\mathrm{bottle}}=879.6\pm 0.6~{}\mathrm{s}$ Pichlmaier et al. (2010); Serebrov et al. (2005); Steyerl et al. (2012); Arzumanov et al. (2015). In contrast, the beam method measures the number of protons produced through $\beta$-decay in a fixed volume from a neutron beam, resulting in a longer lifetime of $\tau_{n}^{\mathrm{beam}}=888.0\pm 2.0~{}\mathrm{s}$ Byrne et al. (1990); Yue et al. (2013). This discrepancy can be reconciled by introducing a new decay channel for the neutron into a light dark baryon $(\chi)$, as $n\rightarrow\chi+\cdots$, where the additional particles ()($\cdots$)( ⋯ ) conserve energy and momentum. Consequently, the branching fraction for this decay mode becomes
@@
[…]
@@ 6 Conclusions
An alternative/additional mechanism of DM accumulation in NSs could be the decay of SM particles into DM. The observed $\sim$8888 s discrepancy between beam and bottle neutron lifetime measurements suggests a possible decay channel involving a dark particle $\chi$ with a branching fraction constrained to be $f_{n\rightarrow\chi}\lesssim 2.7\times 10^{-3}$ (90% CL). In particular, a fermionic DM candidate with mass $m_{\chi}\gtrsim 1.2\,\mathrm{GeV}$ or exhibiting strong self-repulsion could resolve that tension.

### v3

@@ 3.5 Speed of Sound in Dark Matter Admixed Neutron Stars
The stiffness of matter in compact objects is described by the sound speed $c_{s}$, which plays a critical role in preventing a static NS from collapsing into a BH. Unlike the mass or the radius, the sound speed cannot be directly measured by experiments, nor it is easily inferred. However, the speed of sound is still a key parameter for characterizing the NS properties and, in particular, for quantifying how resistant the matter is to compression in extreme conditions. In this regard, Ecker et al. Ecker and Rezzolla (2022) have demonstrated that in light stars (whose $c_{s}$ increases gradually towards the center), the outer layers are soft while the core is rather stiff; whereas heavy NSs have a stiffer outer shell and a softer core (where the sound speed reaches a local minimum).
@@ 5 Neutron Decay to Dark Matter
DM interactions can be constrained by investigating neutron decay within NSs. The longstanding discrepancy in neutron decay lifetime, observed using the beam and bottle methods Berezhiani and Bento (2006); Serebrov et al. (2008); Fornal and Grinstein (2018), provides an avenue to explore these interactions. In the bottle method, neutrons are trapped, making it sensitive to the total neutron width, yielding a slightly shorter lifetime of $\tau_{n}^{\mathrm{bottle}}=879.6\pm 0.6~\mathrm{s}$ Pichlmaier et al. (2010); Serebrov et al. (2005); Steyerl et al. (2012); Arzumanov et al. (2015). In contrast, the beam method measures the number of protons produced through $\beta$-decay in a fixed volume from a neutron beam, resulting in a longer lifetime of $\tau_{n}^{\mathrm{beam}}=888.0\pm 2.0~\mathrm{s}$ Byrne et al. (1990); Yue et al. (2013). This discrepancy can be reconciled by introducing a new decay channel for the neutron into a light dark baryon $(\chi)$, as $n\rightarrow\chi+\cdots$, where the additional particles $(\cdots)$ conserve energy and momentum. Consequently, the branching fraction for this decay mode becomes
@@
A study in Fornal and Grinstein (2018) explores how the $\sim$$8~\mathrm{s}$ discrepancy in neutron lifetime measurements, with a $4\sigma$ tension, could be resolved by introducing an additional neutron decay channel involving a fermionic DM particle $(\chi)$. This particle’s mass falls within the narrow range $m_{p}-m_{e}<m_{\chi}<m_{n}$, ensuring consistency with proton stability. While the simplest decay mode, $n\rightarrow\chi\gamma$, is excluded by the absence of a detected monochromatic photon Tang et al. (2018), alternative invisible decay channels, such as $n\rightarrow\chi\phi$, face stringent constraints from NS observations. The presence of new states within NSs could significantly reduce their maximum mass. Observations of NSs with masses exceeding $0.7~M_{\odot}$ impose strong constraints, requiring $\chi$ particles with a baryon number to have masses greater than $1.2~\mathrm{GeV}$ or to exhibit strongly repulsive self-interactions McKeen et al. (2018); Cline and Cornell (2018). Neutron–mirror neutron oscillations have also been investigated Berezhiani et al. (2018); Berezhiani (2019a, b); Ayres et al. (2022), while an alternative approach, distinct from DM, has been proposed to address the neutron lifetime measurement puzzle, as discussed in Fuwa et al. (2024).
@@ 6 Conclusions
An alternative/additional mechanism of DM accumulation in NSs could be the decay of SM particles into DM. The observed $\sim$88 s discrepancy between beam and bottle neutron lifetime measurements suggests a possible decay channel involving a dark particle $\chi$ with a branching fraction constrained to be $f_{n\rightarrow\chi}\lesssim 2.7\times 10^{-3}$ (90% CL). In particular, a fermionic DM candidate with mass $m_{\chi}\gtrsim 1.2\,\mathrm{GeV}$ or exhibiting strong self-repulsion could resolve that tension.